OpenStudy (anonymous):

help! write the equation of the line with slope: 6/7 y intercept: -9/2

OpenStudy (anonymous):

To solve this problem, we need to understand how the equation of the line works: Y = mx + b m = slope b = y intercept. So, replacing m with the slope value and b with y intercept we get: y = (6/7)x + (-9/2) which is: y = 6/7x - 9/2